OpenStudy (anonymous):

evaluate the function f(x)=4•7^x for x=-1 and x=2 . show your work.

OpenStudy (bibby):

just plug in the values and evaluate. recall that \(\large x^{-n}=\dfrac{1}{x^n}\)

OpenStudy (anonymous):

is that the answer? i dont know how to do this.. I've been stuck on this question for days

OpenStudy (bibby):

is what the answer? the answer is 2 different numbers. plug in x= 2 and evaluate

OpenStudy (anonymous):

I'm confused!! I really can't do this crap. if I did that then it'd be 2^-n=1/2^n . how the hell am I supposed to evaluate that?

OpenStudy (bibby):

\(\huge f(x)=4*7^x\\x=2\\\large f(2)=4*7^2\)

OpenStudy (bibby):

we're trying to find f(-1) and f(2)

OpenStudy (anonymous):

okay but how do you do that

OpenStudy (bibby):

use a calculator? have you not learned exponents yet?

OpenStudy (bibby):

the number up on top represents how much times it multiplies itself. for example \(3^2=3*3\\6^3=6*6*6 \)

OpenStudy (anonymous):

I don't have a calculator or else I would. & yeah I have. but all these variables and cramp confuse me. obviously I need help or I wouldn't have asked the question

OpenStudy (anonymous):

ok

OpenStudy (bibby):

so just expand and multiply by hand

OpenStudy (bibby):

chill out with the attitude yo, I have homework I'm not doing right now, be grateful

OpenStudy (bibby):

what is 7^2?